OpenStudy (anonymous):

square root of xy divided by the 5th root of 32xy. simplify the expression and express the answer using rational exponents. Assume that x and y denote positive numbers.

OpenStudy (anonymous):

\[\frac{\sqrt{xy}}{\sqrt[5]{32xy}}=2^{-1}x^\frac{3}{10}y^\frac{3}{10}\]